OpenStudy (anonymous):

Rationalize the denominator and simplify. ^3sqrt(4/9) I know I'm wrong but I think its ^3sqrt12/3

OpenStudy (anonymous):

i think you are correct

OpenStudy (anonymous):

you have \[\sqrt[3]\frac{4}{9}\times \sqrt[3]\frac{3}{3}=\frac{\sqrt[3]{12}}{3}\]

OpenStudy (anonymous):

Yes